OpenStudy (anonymous):

How do i solve ((a-2)^2)-4

OpenStudy (aayushi.somani):

[(a-2)^2]-4 =a^2+4-4a-4 =a^2-4a

OpenStudy (amistre64):

a solution depends on what it is your trying to accomplish .... otherwise its the locus of all the points of \(a\in R\)

OpenStudy (anonymous):

oh i mean factorizing

OpenStudy (amistre64):

hmm, then the first poster has done a few steps ... another idea is to apply a different of squares

OpenStudy (amistre64):

(x^2 - y^2) = (x + y) ( x - y)

OpenStudy (anonymous):

oh yeah i remember that

OpenStudy (amistre64):

im surprised that i do :)

OpenStudy (amistre64):

\[(a-2-2)~(a-2+2)\]

OpenStudy (anonymous):

:D thanks

OpenStudy (amistre64):

good luck